OpenStudy (firejay5):

What would be an equation in y - intercept form be for points: (4,3) and (7, -2) graphed

OpenStudy (firejay5):

those points are graphed and I have the slope; which is - 5/3

OpenStudy (firejay5):

In order to find the slope: y2 - y1 divided by x2 - x1

OpenStudy (anonymous):

When you have slope, plug m = -5/3 and any point coordinate into Point slope form

OpenStudy (firejay5):

how would it work out then

OpenStudy (anonymous):

Do you know Point-Slope form?

OpenStudy (firejay5):

I don't think show me

OpenStudy (anonymous):

y - y1 = m ( x - x1)

OpenStudy (firejay5):

How would it work in an example or problem working on now

OpenStudy (anonymous):

Choose the point (4, 3) then plug its value, also slope into the formula

OpenStudy (firejay5):

Show me how's it's plugged in

OpenStudy (firejay5):

@Chlorophyll it's too confusing

OpenStudy (anonymous):

Why don't you plug m and x1 in?

OpenStudy (firejay5):

is there another way of doing it

OpenStudy (anonymous):

NO, there's only 1 and 1 way! Believe me if there's another easier way I already showed you!

OpenStudy (anonymous):

If you can't follow my instruction, I'll give up with whatever your posts later !

OpenStudy (firejay5):

I have a graphing calculator and is trying to figure out in the table what equation goes with (4,3) and (7,-2) and I already have y = - 5/3x ........

OpenStudy (anonymous):

Why would you rather to keep complaining "confused" in stead of plug the numbers in???

OpenStudy (anonymous):

I'll return one more time to check it!

OpenStudy (firejay5):

see I have y - 3 = - 5/3x ...... should I distrubute

OpenStudy (firejay5):

before that I had y -3 = - 5/3 x - (4) <---- what do I do here

OpenStudy (firejay5):

okay it's the same

OpenStudy (firejay5):

I got - 20/3

OpenStudy (firejay5):

I am probably wrong, but all I want is the answer that's all the teacher cares

OpenStudy (firejay5):

plus a little work shown

OpenStudy (firejay5):

that's what I got when I distributed

OpenStudy (firejay5):

y = - 5/3x - 20/3

OpenStudy (firejay5):

0 idk, making me figure this out isn't helping me

OpenStudy (firejay5):

okay

OpenStudy (firejay5):

you're smarter then me

OpenStudy (firejay5):

It would be nice to tell me the answer if I have 1/2 of it: y = -5/3 x + ?

OpenStudy (firejay5):

Is it 5, 4, or something. Well post it on here. Me logging off

OpenStudy (anonymous):

I don't know if you still need help, but I can try to explain it. (y-y1)=m(x-x1) is the point slope form where m is the slope of the line y1 is the y coordinate of a point on that line and x1 is the x coordinate of the same point you just plug in the numbers and solve for y as chlorophyll did above this form is very useful when you only have two points to find an equation

OpenStudy (firejay5):

Guys here's the point slope m = \[\left[\begin{matrix} y2 - y1 \\ x2 - x2\end{matrix}\right]\]

OpenStudy (firejay5):

It's in a fraction meaning dividing

OpenStudy (firejay5):

I don't need it, because @Chlorophyll WAS CONFUSING ME

OpenStudy (anonymous):

Okie, so let start from the beginning :)